OpenStudy (anonymous):

I have finished an algebra review, and I was wondering if anyone would be willing to check my answers? (:

OpenStudy (anonymous):

Hmm I can :D

OpenStudy (anonymous):

1. Solve, using linear combination. 3x – y = 4 2x + y = 6 A. (–1, –7) B. (3, 0) C. (0, –4) D. (2, 2) I believe that this one is D.

OpenStudy (anonymous):

lol!!

OpenStudy (anonymous):

2. Use the graph to state the solution for the system. x + 2y = 4 (line a) 3x – 2y = 4 (line b) A. (–2, –1) B. (2, 1) C. (–2, 1) D. (2, –1) I think that that number two is B.

OpenStudy (anonymous):

3. Solve this system of equations using any method. 2x – y = 5 3x + 2y = 4 A. (5, 5) B. (0, 2) C. (2, –1) D. (–1, –7) I think that this one is C

OpenStudy (anonymous):

Which is a system of equations to solve the following problem? Let c be the number of child tickets and a be the number of adult tickets. Each child ticket for a ride costs $2, while each adult ticket costs $6. If the ride collected a total of $148, and 38 tickets were sold, how many of each type of ticket were sold? Here is my work: 2c+6a=148 a+c=38

OpenStudy (anonymous):

Each child ticket for a ride costs $2, while each adult ticket costs $6. If the ride collected a total of $148, and 38 tickets were sold, how many of each type of ticket were sold? A. 19 child and 19 adult B. 14 child and 20 adult C. 148 adult and 38 child D. 20 child and 18 adult This one is B

OpenStudy (jim766):

1st 2 look good

OpenStudy (anonymous):

In a recent golf match, Becky’s score was 5 less than Cathy’s score. Their combined scores totaled 185. Which of the following pairs of equations can be used to determine their scores? Let b represent Becky’s score and c represent Cathy’s score. A. b + c = 185 b – c = 5 B. c = b – 5 b + c = 185 C. b = c – 5 b + c = 185 D. b + c = 5 185 + b = c This one is A

OpenStudy (anonymous):

Solve the following system of equations. d + e = 6 d – e = 4 I don't think that there is a solution to this equation.

OpenStudy (jim766):

look back at the ticket problem you said was b

OpenStudy (anonymous):

Which is the graph of the system of equations? 2x + 3y = 12 4x + y = 8 I believe that this is the correct graph:

rishavraj (rishavraj):

d + e = 6 & d - e = 4.... theres a solution for tht

OpenStudy (anonymous):

What would be the solution then?

rishavraj (rishavraj):

d = 5 & e = 1

OpenStudy (jim766):

add them straight down

OpenStudy (anonymous):

Solve the following system of equations by linear combination: 3x + 2y = 7 5x – 4y = 19 I got (3,-1)

OpenStudy (anonymous):

Those are all of the questions, and all of my answers. Can you guys please tell me how I did, if you know? :)

OpenStudy (jim766):

last one looks good

rishavraj (rishavraj):

(3, -1) is correct...

OpenStudy (jim766):

I get (20,18) for that ticket problem above

rishavraj (rishavraj):

question on golf match...ur answer is wrong...:(

OpenStudy (anonymous):

What about the others? :)

rishavraj (rishavraj):

Ticket one answer is also wrong:(

rishavraj (rishavraj):

other r correct except the two mentiond above

OpenStudy (anonymous):

Okay thank you!

rishavraj (rishavraj):

u r most welcome....anytime ;)